# Robust Hyperbolic Learning with Curvature-Aware Optimization

### v3

@@ 4.2 EEG Classification Problem
We compare our model HyperMAtt on three common EEG datasets for the domain of EEG classification. These datasets are MI (Motor Imagery) (Brunner et al., 2008), SSVEP (Steady-State Visual Evoked Potentials) (Nikolopoulos, 2021) and ERN (Error-Related Negativity) (Margaux et al., 2012) representing three different tasks, motor imagery, visual stimuli, and error recognition, respectively. Here, we follow the same preprocessing procedure and train/val/test splits as MAtt. For the datasets MI and SSVEP, we use Accuracy as an evaluation metric and AUC.
For the application domain of EEG, we evaluate our method against current state-of-the-art models for EEG Classification including Inception (Burchert et al., 2024), MAtt (Pan et al., 2022), MBEEGSE (Altuwaijri et al., 2022), FBCNet (Mane et al., 2021), TCNet-Fusion (Musallam et al., 2021), EEG-TCNet (Ingolfsson et al., 2020), SCCNet (Wei et al., 2019), EEGNet (Lawhern et al., 2018), and ShallowConvNet (Schirrmeister et al., 2017).
@@ Results
In Table 2 we show the performance of HyperMAtt compared to the current state-of-the-art baselines, where the results were aggregated from (Pan et al., 2022; Burchert et al., 2024). The performance for HyperMAtt is measured across 3 runs and we report the mean and standard deviation. Here, our new optimizer archives state-of-the-art results for SSVEP and ERN, the two comparatively smaller datasets, which are more prone to overfitting. We also show the improved performance of our RAdamW optimizer vs the existing RAdam by training the model using both and comparing the results in Table 8. We also achieve these results in significantly less time/epoch compared to MAtt. Specifically for MI 0.077s/epoch vs 3.4s/epoch, for SSVEP 0.081s/epoch vs 4s/epochs, and for ERN 0.061s/epoch vs 1.9s/epoch.
Models MI SSVEP ERN ShallowConvNet 61.84$\pm$6.39 56.93$\pm$6.97 71.86$\pm$2.64 EEGNet 57.43$\pm$6.25 53.72$\pm$7.23 74.28$\pm$2.47 SCCNet 71.95$\pm$5.05 62.11$\pm$7.70 70.93$\pm$2.31 EEG-TCNet 67.09$\pm$4.66 55.45$\pm$7.66 77.05$\pm$2.46 TCNet-Fusion 56.52$\pm$3.07 45.00$\pm$6.45 70.46$\pm$2.94 FBCNet 71.45$\pm$4.45 53.09$\pm$5.67 60.47$\pm$3.06 MBEEGSE 64.58$\pm$6.07 56.45$\pm$7.27 75.46$\pm$2.34 Inception 62.85$\pm$3.21 62.71$\pm$2.95 73.55$\pm$5.08 MAtt 74.71$\pm$5.01 65.50$\pm$8.20 76.01$\pm$2.28 HyperMAtt (ours) 74.12$\pm$2.91 68.10$\pm$2.41 78.01$\pm$1.30 Increase in % -0.79 3.97 1.25
4.3 Standard Image Classification Problem
Problem Setting and Reference Model
In their work, Bdeir et al. (2024) proposed a fully hyperbolic 2D convolutional layer by breaking down the convolution operation into a window-unfolding step followed by a modified version of the Lorentz Linear Layer from Chen et al. (2022). This approach ensured that the convolution outputs remained on the hyperboloid. However, the manual patch creation combined with matrix multiplication made the computation extremely expensive, as it prevented the use of highly optimized CUDA implementations for convolutions. As such, the authors proposed a two versions of their hyperbolic ResNet classifier. A fully hyperbolic model with all Lorentz ResNet blocks (HCNN) and a hybrid encoder model with alternating Euclidean and hyperbolic blocks (HECNN).
Reference Optimization
To adapt our methods for both models, we introduce the distance rescaling function after every convolution in the encoder. Additionally, we attempt to improve the performance by performing a parametrization trick on existing CUDA convolutions to ensure hyperbolic outputs. This is done by parametrizing the convolution weight as a rotation operation before passing the input, and then applying a boost operation afterwards. The end result is then equivalent to the convolution operation by Bdeir et al. (2024) while mitigating computation complexity. Specific implementation details can be found in Appendix B. We denote our Hybrid and Fully hyperbolic models as HECNN+ and HCNN+ respectively.
Experimental Goals
In their paper, (Bdeir et al., 2024) notice lower performance in the fully hyperbolic models compared to the hybrid models and attribute this to instability in training. Additionally, they cite crashes and model divergence when learning the curvature instead of setting it to constant. This would then be an ideal setting to test the optimization schema and rescaling function and their impact on performance and stability. We do not use the RAdamW optimizer in this setting as RSGD is already used and employs proper L2-regularization.
Experimental Setup
We follow the experimental setup and hyperparameters of Bdeir et al. (2024) and rely on three main datasets: CIFAR10, CIFAR100, and Mini-Imagenet. For ResNet-18, we employ HECNN+, following the structure outlined in Bdeir et al. (2024) with alternating Euclidean and hyperbolic blocks, and integrate our efficient convolution operation. We decouple the encoder and decoder manifolds, allowing each to independently learn its own curvature for enhanced model flexibility. We also implement ResNet-50 models of the proposed hyperbolic hybrid architectures using similarly constructed bottleneck blocks.
Results
For the ResNet-18 experiments, Table 3 shows that the new models are able to remain stable while learning the curvature. Additionally, we see significant performance improvements between HCNN and HCNN+, where our proposed architecture now matches the performance of the hybrid encoders. We hypothesize that the improved scaling function helps mitigate the previous performance inconsistencies. However, the performance difference between the hybrid models is not significant, we hypothesize this is due to the alternating architecture which could limit the effect and instability of hyperbolic components. Finally, both models manage to maintain performance while reducing the memory footprint and runtime by approximately $25-50\%$ and $18-25\%$ respectively. One other benefit that we find from learning the curvature is quicker convergence where the model is able to reach convergence in 130 epochs vs the 200 epochs required by a static curve model.
Table 3: Runtime Analysis for the ResNet-18 models. We report classification accuracy (%) and estimate the mean and standard deviation from five runs. The best performance is highlighted in bold (higher is better).
CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.23)$ For Cifar100 $(\delta_{rel}=0.20)$ Euclidean $77.72_{\pm 0.15}$ 1.2GB - 12s $65.19_{\pm 0.12}$ Hybrid ($\mathbb{P}$) $77.19_{\pm 0.50}$ $64.93_{\pm 0.38}$ Hybrid ($\mathbb{L}$) ${78.03_{\pm 0.21}}$ ${65.63_{\pm 0.10}}$ ResNet ($\mathbb{P}$) $76.60_{\pm 0.32}$ $62.01_{\pm 0.56}$ HECNN ($\mathbb{L}$) ${78.76_{\pm 0.24}}$ 4.3GB - 100s ${65.96_{\pm 0.18}}$ HECNN+ ${78.80_{\pm 0.12}}$ 3GB - 80s ${65.98_{\pm 0.11}}$ HCNN ($\mathbb{L}$) ${78.07_{\pm 0.17}}$ 10GB - 175s ${65.71_{\pm 0.13}}$ HCNN+ $\bm{78.81_{\pm 0.19}}$ 5GB - 140s $\bm{66.12_{\pm 0.14}}$
In the ResNet-50 experiments in Table 4, HECNN+ significantly outperforms both the Euclidean model and the base hybrid model across both datasets. We also see a $\sim 48\%$ reduction in memory usage and $\sim 66\%$ reduction in runtime. We attribute this improvement to efficient closed-source convolution operations we can now leverage.
Table 4: Runtime Analysis for ResNet-50 models. We report classification accuracy (%) and the best performance is highlighted in bold (higher is better).
CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ Euclid $78.52$ 4.5GB - 30s $66.23$ HECNN 79.8379.83{79.83}79.83 15.6GB - 300s $66.30$ HECNN+ $\bm{80.86}$ 8.1GB - 100s $\bm{67.18}$
4.4 VAE Image Generation
Problem Setting and Reference Model
Previous research has highlighted the ability of hyperbolic neural networks to perform better when using lower dimensional embeddings due to the more expressive space (Nagano et al., 2019; Mathieu et al., 2019; Ovinnikov, 2019; Hsu et al., 2020). A natural extension to this would be implementing the models for VAEs which rely on smaller latent embedding dimensions to encode the inputs. Bdeir et al. (2024) perform this study on image generation and reconstruction for their fully hyperbolic models.
Experimental Setup
We reproduce the experimental setup from (Bdeir et al., 2024) and re-implement the fully hyperbolic VAE using our new efficient convolution and transpose convolution layers. We also use curvature learning with our adjusted Riemannian SGD learning scheme.
Results
@@
Our fully hyperbolic VAE implementation outperforms on both datasets, as shown in Table 5, while using 2.5x less memory and training 3x faster. This highlights the effectiveness of our curvature learning process and efficient model components.
@@ 5 Conclusion
In this work, we propose a robust curvature-aware optimization framework for hyperbolic deep learning, addressing challenges in stability, computational efficiency, and overfitting. By introducing Riemannian AdamW, a novel distance rescaling function, and leveraging efficient CUDA implementations, our approach achieved performance improvements in hierarchical metric learning, EEG classification, and image classification tasks, while reducing computational costs, highlighting the practicality of our methods for large-scale applications. These findings emphasize the importance of proper curvature adaptation in hyperbolic learning, paving the way for future research in optimizing hyperbolic embeddings across diverse fields.
6 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. We are committed to contributing positively to society and human well-being, while respecting the privacy of the subjects involved in our evaluation. Our experiments utilize three established EEG datasets: MI, SSVEP, and ERN, which involve data from human subjects. These datasets are fully anonymized and do not contain any personally identifiable information. Additionally, they are publicly available and widely used within the machine learning community.
@@ Appendix A Operations in hyperbolic geometry
A parallel transport operation $\mathrm{PT}^{K}_{{\bm{x}}\rightarrow{\bm{y}}}\left({\bm{v}}\right)$ describes the mapping of a vector on the manifold $\bm{v}$ from the tangent space of $\bm{x}\in\mathbb{L}$ to the tangent space of $\bm{y}\in\mathbb{L}$. This operation is given as $\mathrm{PT}^{K}_{{\bm{x}}\rightarrow{\bm{y}}}\left({\bm{v}}\right)=\bm{v}+\frac{\langle\bm{y},\bm{v}\rangle_{\mathbb{L}}}{K-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}(\bm{x}+\bm{y}).$
Lorentzian Centroid (Law et al., 2019)
Also denoted as $\bm{\mu}_{\mathbb{L}}$, is the weighted centroid between points on the manifold based on the Lorentzian square distance. Given the weights $\bm{\nu}$, $\bm{\mu}=\frac{\sum_{i=1}^{m}\nu_{i}\bm{x}_{i}}{\sqrt{1/K}\left|||\sum_{i=1}^{m}\nu_{i}\bm{x}_{i}||_{\mathcal{L}}\right|}.$
@@
In the Lorentz model, linear transformations preserving the structure of spacetime are termed Lorentz transformations. A matrix $\mathrm{\textbf{A}}^{(n+1)\times(n+1)}$ is defined as a Lorentz transformation if it provides a linear mapping from $\mathbb{R}^{n+1}$ to $\mathbb{R}^{n+1}$ that preserves the inner product, i.e., $\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{y}\rangle_{\mathbb{L}}=\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$ for all $\bm{x},\bm{y}\in\mathbb{R}^{n+1}$. The collection of these matrices forms an orthogonal group, denoted $\bm{O}(1,n)$, which is commonly referred to as the Lorentz group.
@@
$\bm{O}^{+}(1,n)={\mathrm{\textbf{A}}\in\mathbb{R}^{(n+1)\times(n+1)}|\forall\bm{x}\in\mathbb{L}^{n}_{K}:\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{x}\rangle_{\mathbb{L}}=-\frac{1}{K},(\mathrm{\textbf{A}}\bm{x})_{0}>0)}.$ (6)
Each Lorentz transformation can be decomposed via polar decomposition into a Lorentz rotation and a Lorentz boost, expressed as $\mathrm{\textbf{A}}=\mathrm{\textbf{R}}\mathrm{\textbf{B}}$ (Moretti, 2002). The rotation matrix R is designed to rotate points around the time axis and is defined as
@@
In contrast, the Lorentz boost applies shifts along spatial axes given a velocity vector $\bm{v}\in\mathbb{R}^{n}$ with $||\bm{v}||<1$, without altering the time axis.
@@
with $\gamma=\frac{1}{\sqrt{1-||\bm{v}||^{2}}}$. However, this can also be any operation that scales the norms of the space values without changing the vector orientation.
@@ Appendix B Convolution Trick
To address the computational requirements issue, we adopt an alternative definition of the Lorentz Linear layer from Dai et al. (2021), which decomposes the transformation into a Lorentz boost and a Lorentz rotation. Using this definition, we replace the matrix multiplication employed by Bdeir et al. (2024) for the spatial dimensions and time component projection with a learned rotation operation and a Lorentz boost. Additionally, we can achieve the rotation operation using a parameterization of the convolution weights while still relying on the CUDA convolution implementations, significantly improving computational efficiency.
@@
$\bm{out}=\text{LorentzBoost(DistanceRescaling(RotationConvolution(}\bm{x}\textbf{)))}$)) (9)
@@
Algorithm 2 Lorentz Convolution Parametarization
  $\bm{W}\in\mathbb{R}^{C_{in},C_{out},K_{width},K_{length}}$
  function Adaptweight
     if $K_{width}\cdot K_{length}\cdot C_{in}>C_{out}$ then
        return $\bm{W}$
     else
        $\bm{W}=\bm{W}.reshape(K_{width}\cdot K_{length}\cdot C_{in},C_{out})$
        $\bm{\hat{W}}_{core}=$ Transform ($\bm{W}$)
        return $\bm{\hat{W}}.reshape(C_{in},C_{out},K_{width},K_{length})$
     end if
  end function
@@ Appendix C Scaling Lorentzian Vectors
Appendix D Ablations
D.1 Image Classification
We test the effect of individual model components in Table 6. Each subsequent model involves the default architecture presented in the experimental setup minus the mentioned component. As we can see, the best results are achieved when all the architectural components are included. In the case of attempting to learn the curvature without the proposed optimizer schema, the model breaks completely down due to excessive numerical inaccuracies.
Table 6: Ablation experiments for Resnet-50 models on Cifar-100.

### v4

@@ 4.4 VAE Image Generation
We reproduce the experimental setup from [3] and re-implement the fully hyperbolic VAE using our new efficient convolution and transpose convolution layers. We also use curvature learning with our adjusted Riemannian SGD learning scheme.
@@ Results
Our fully hyperbolic VAE implementation outperforms on both datasets, as shown in Table˜4, while using 2.5x less memory and training 3x faster. This highlights the effectiveness of our curvature learning process and efficient model components.
@@ 5 Conclusion
In this work, we propose a robust curvature-aware optimization framework for hyperbolic deep learning, addressing challenges in stability, computational efficiency, and overfitting. By introducing Riemannian AdamW, a novel distance rescaling function, and leveraging efficient CUDA implementations, our approach achieved performance improvements in hierarchical metric learning, EEG classification, and image classification tasks, while reducing computational costs, highlighting the practicality of our methods for large-scale applications. These findings emphasize the importance of proper curvature adaptation in hyperbolic learning, paving the way for future research in optimizing hyperbolic embeddings across diverse fields. Further work should be done however to address the limitations on theoretical reasoning and efficiency presented in Appendix F.
6 Acknowledgements
Ahmad Bdeir was funded by the European Union’s Horizon 2020 research and innovation programme under the SustInAfrica grant agreement No 861924, as well as Horizon Europe Project BioMonitor4CAP grant agreement No 101081964.
@@ Appendix A Operations in hyperbolic geometry
A parallel transport operation $\mathrm{PT}^{K}_{{\bm{x}}\rightarrow{\bm{y}}}\left({\bm{v}}\right)$ describes the mapping of a vector on the manifold $\bm{v}$ from the tangent space of $\bm{x}\in\mathbb{L}$ to the tangent space of $\bm{y}\in\mathbb{L}$. This operation is given as $\mathrm{PT}^{K}_{{\bm{x}}\rightarrow{\bm{y}}}\left({\bm{v}}\right)=\bm{v}+\frac{\langle\bm{y},\bm{v}\rangle_{\mathcal{L}}}{K-\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}}(\bm{x}+\bm{y}).$
Lorentzian Centroid [22]
[…]
Optimization Operations
egrad2rgrad converts a Euclidean gradient (computed in the ambient space) to a Riemannian gradient by projecting it onto the tangent of the corresponding hyperbolic parameter. For the Lorentz manifold this is then:
$\text{egrad2rgrad}(\mathbf{x},v)=v+\frac{\langle v,\mathbf{x}\rangle_{\mathcal{L}}}{K}\mathbf{x}.$
The retraction maps the updated hyperbolic vector from the tangent space back onto the manifold. In the case of the hyperboloid this is the proposed exponential mapping equation.
@@
In the Lorentz model, linear transformations preserving the structure of spacetime are termed Lorentz transformations. A matrix $\mathrm{\textbf{A}}^{(n+1)\times(n+1)}$ is defined as a Lorentz transformation if it provides a linear mapping from $\mathbb{R}^{n+1}$ to $\mathbb{R}^{n+1}$ that preserves the inner product, i.e., $\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{y}\rangle_{\mathcal{L}}=\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}$ for all $\bm{x},\bm{y}\in\mathbb{R}^{n+1}$. The collection of these matrices forms an orthogonal group, denoted $\bm{O}(1,n)$, which is commonly referred to as the Lorentz group.
@@
$\bm{O}^{+}(1,n)={\mathrm{\textbf{A}}\in\mathbb{R}^{(n+1)\times(n+1)}|\forall\bm{x}\in\mathbb{L}^{n}_{K}:\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{x}\rangle_{\mathcal{L}}=-\frac{1}{K},(\mathrm{\textbf{A}}\bm{x})_{0}>0)}.$ (4)
Each Lorentz transformation can be decomposed via polar decomposition into a Lorentz rotation and a Lorentz boost, expressed as $\mathrm{\textbf{A}}=\mathrm{\textbf{R}}\mathrm{\textbf{B}}$ [30]. The rotation matrix R is designed to rotate points around the time axis and is defined as
@@
In contrast, the Lorentz boost applies shifts along spatial axes given a velocity vector $\bm{v}\in\mathbb{R}^{n}$ with $\bm{v}||<1$, without altering the time axis.
@@
[…]
(a) Tangent space at a random point on the manifold
(b) Tangent space at the origin
Figure 1: Tangent planes of a hyperboloid with curvature -1 relative to another hyperboloid with curvature -0.7. Tangential properties between manifolds are better respected at the origin where tangents remain parallel.
@@ Appendix B Convolution Trick
To address the computational requirements issue, we adopt an alternative definition of the Lorentz Linear layer from Dai et al. [8], which decomposes the transformation into a Lorentz boost and a Lorentz rotation. Using this definition, we replace the matrix multiplication employed by Bdeir et al. [3] for the spatial dimensions and time component projection with a learned rotation operation and a Lorentz boost. Additionally, we can achieve the rotation operation using a parameterization of the convolution weights while still relying on the CUDA convolution implementations, significantly improving computational efficiency.
@@
$\bm{out}=\text{LorentzBoost(DistanceRescaling(RotationConvolution(}\bm{x}\textbf{)))}$ (7)
@@
Algorithm 3 Lorentz Convolution Parameterization
[…]
2:function AdaptWeight
3:  if $K_{\text{width}}\cdot K_{\text{length}}\cdot C_{\text{in}}\leq C_{\text{out}}$ then
4:   $\bm{W}\leftarrow\text{reshape}(\bm{W},K_{\text{width}}\cdot K_{\text{length}}\cdot C_{\text{in}},C_{\text{out}})$
5:   $\bm{\hat{W}}_{\text{core}}\leftarrow\text{Transform}(\bm{W})$
6:   $\bm{W}\leftarrow\text{reshape}(\bm{\hat{W}},C_{\text{in}},C_{\text{out}},K_{\text{width}},K_{\text{length}})$
7:  end if
8:  return $\bm{W}$
9:end function
@@ Appendix C Scaling Lorentzian Vectors
Appendix D Proofs
In the following, we show that the distances and angles between the origin and the hyperbolic points are preserved. This is relatively trivial but we add it for sake of clarity. For this discussion, we work in the Lorentz manifold and repeat the definitions here for easier referencing.
$\mathcal{L}^{n}:=\{\mathbf{x}=[x_{t},\mathbf{x_{s}}]\in\mathbb{R}^{n+1}~|~\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=-K,~x_{t}>0\}$, $\mathbf{x},\mathbf{y}>_{\mathcal{L}}=-x_{t}y_{t}+\mathbf{x}_{s}^{T}\mathbf{y}_{s}$ and the origin is $O=[O_{0},\mathbf{O}_{s}]=[\sqrt{K},0,\dots,0]$.
We also define two hyperboloids $H$ and $H^{\prime}$ with curvatures K-{K} and $K^{\prime}$} respectively. Let $x,y$ be two points on $H$ and $v,w\in T_{O}H$ be their logarithmic map onto the tangent space at the origin of $H$.
Proof of distance preservation
We can show that distances to the origin are preserved because $d_{\mathbb{L}}(\mathbf{x},\mathbf{O})$ collapses to $\lvert\lvert v\rvert\rvert_{euclid}$. This gives $d_{\mathbb{L}}(\mathbf{x},\mathbf{O})=\lvert\lvert v\rvert\rvert_{euclid}=d_{\mathbb{L}}(\mathbf{x^{\prime}},\mathbf{O^{\prime}})$.
Proof of angle preservation
We can similarly show that angles w.r.t to the origin are preserved. The angle $\theta$ between $x$ and $y$ at $O$ is actually computed via the Euclidean inner product in $T_{O}H$ using
$\cos\theta\frac{<v,w>}{\lvert\lvert v\rvert\rvert\lvert\lvert w\rvert\rvert}$
As such the angle w.r.t. origins also does not change when moving between manifolds of different curvature using this method since $v$ and $w$ don’t change.
Appendix E Additional Experiments
E.1 Image Classification
Problem Setting and Reference Model
In their work, Bdeir et al. [3] proposed a fully hyperbolic 2D convolutional layer by breaking down the convolution operation into a window-unfolding step followed by a modified version of the Lorentz Linear Layer from Chen et al. [6]. This approach ensured that the convolution outputs remained on the hyperboloid. However, the manual patch creation combined with matrix multiplication made the computation extremely expensive, as it prevented the use of highly optimized CUDA implementations for convolutions. As such, the authors proposed a two versions of their hyperbolic ResNet classifier. A fully hyperbolic model with all Lorentz ResNet blocks (HCNN) and a hybrid encoder model with alternating Euclidean and hyperbolic blocks (HECNN).
Table 5: Performance and runtime Analysis for the ResNet-18 models. We report classification accuracy (%) and estimate the mean and standard deviation from five runs. The best performance is highlighted in bold (higher is better).
CIFAR-100 Tiny-ImageNet VRAM tepoch
$(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ For Cifar100 For Cifar100
Euclidean $77.72_{\pm 0.15}$ $65.19_{\pm 0.12}$ 1.2GB 12s
Hybrid Poincaré [16] $77.19_{\pm 0.50}$ $64.93_{\pm 0.38}$ - -
Hybrid Lorentz [3] $78.03_{\pm 0.21}$} $65.63_{\pm 0.10}$} - -
Poincaré ResNet [40] $76.60_{\pm 0.32}$ $62.01_{\pm 0.56}$ - -
HECNN [3]($\mathbb{L}$) $78.76_{\pm 0.24}$} $65.96_{\pm 0.18}$} 4.3GB 100s
HECNN+ (ours) $78.80_{\pm 0.12}$} $65.98_{\pm 0.11}$} 3GB 80s
HCNN [3] ($\mathbb{L}$) $78.07_{\pm 0.17}$} $65.71_{\pm 0.13}$} 10GB 175s
HCNN+ (ours) $\bm{78.81_{\pm 0.19}}$ $\bm{66.12_{\pm 0.14}}$ 5GB 140s
Resnet-18 Experiments
Table˜5 shows that the new models are able to remain stable while learning the curvature. Additionally, we see significant performance improvements between HCNN and HCNN+, where our proposed architecture now matches the performance of the hybrid encoders. We hypothesize that the improved scaling function helps mitigate the previous performance inconsistencies. However, the performance difference between the hybrid models is not significant, we hypothesize this is due to the alternating architecture which could limit the effect and instability of hyperbolic components. Finally, both models manage to maintain performance while reducing the memory footprint and runtime by approximately $25-50\%$ and $18-25\%$ respectively.
Table 6: Resnet-50 Ablations on Cifar-100.